OpenStudy (anonymous):

is anybody good with perimeter and area? because i might need some help. the questions are gonna be in the comment

OpenStudy (anonymous):

I am good what excatly is ur question?

OpenStudy (texaschic101):

* sitting here eating my cereal...patient;y waiting :)

OpenStudy (anonymous):

whats the ?

OpenStudy (anonymous):

ill type the questions in a sec. just give me a little sec please

OpenStudy (anonymous):

Cereal at this time of day?

OpenStudy (anonymous):

I eat cereal at midnight psshh

OpenStudy (anonymous):

I can eat ceral for breakfast, lunch & dinner

OpenStudy (anonymous):

Which word correctly describes the hexagon? A.equiangular B.equilateral C.regular

OpenStudy (anonymous):

a. equiangular

OpenStudy (anonymous):

a

hartnn (hartnn):

all sides equal = equilateral

OpenStudy (anonymous):

Those angles are not all equal, but the sides are.

OpenStudy (anonymous):

Thanx for becoming my fan Awesomenss 3291

OpenStudy (anonymous):

do yall mind helping answering a bounce of questions for me

hartnn (hartnn):

all angles equal = equiangular both all sides and all angles equal = regular

OpenStudy (anonymous):

u welcome

OpenStudy (anonymous):

hartnn is right

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

just give me a little bit and i will have another question

OpenStudy (anonymous):

Well I'll sit here and eat my Pop-Tarts.

OpenStudy (anonymous):

What is the approximate circumference of a circle with a radius of 40 inches? Use π ≈ 3.14.

OpenStudy (anonymous):

i love poptarts

OpenStudy (anonymous):

\[C=2πr \]

OpenStudy (anonymous):

when im done can somebody give me a medal?

OpenStudy (anonymous):

ok thanks. i will have a few more questions

OpenStudy (anonymous):

Can I have a medal?

OpenStudy (anonymous):

2*3.14*40=251.2

OpenStudy (anonymous):

thanks i got the answer right before you typed it

OpenStudy (anonymous):

Are there anymore questions?

OpenStudy (anonymous):

ya just give me one sec

OpenStudy (anonymous):

Okay I m waiting, just eating some toast and eggs and pancakes

OpenStudy (anonymous):

Breakfast in the evening jokes.

OpenStudy (anonymous):

I forgot to say bacon.

OpenStudy (anonymous):

wow do you just love breakfast foods or something

OpenStudy (anonymous):

Yes I do.

OpenStudy (anonymous):

A circular pan has a diameter of 8 inches. What is the exact area of the pan? A.8π square inches B.16π square inches C.256π square inches D.64π square inches

OpenStudy (anonymous):

Well, texaschic101 started it.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

The answer is 50.27

OpenStudy (anonymous):

\[A=πr^{2}\]

OpenStudy (anonymous):

3.14*16

OpenStudy (anonymous):

Sorry wrong answer

OpenStudy (anonymous):

So, the answer is B.

OpenStudy (anonymous):

Are there anymore questions? Thanx for being my gan OrangeMaster

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

ok great thanks hope i get it right. and ya. are y'all just in the mood to answer questions or what?

OpenStudy (anonymous):

I like answering questions.

OpenStudy (anonymous):

No problem, and I do not have anything else to do.

OpenStudy (anonymous):

...and eating steak for breakfast

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

awesome. and i love steak too.

OpenStudy (anonymous):

i have a question what is the area of the figure? A.240 in2 B.126 in2 C.360 in2 D.282 in2

OpenStudy (anonymous):



OpenStudy (anonymous):

U break it into different parts, 7*7 + 12*20 = 289

OpenStudy (anonymous):



OpenStudy (anonymous):

12*20=240 7*6=42 240+42=282

OpenStudy (anonymous):

Sorry Orange Master is right 282